OpenStudy (anonymous):

2x+7y=-1 4x-3y=-19 what is the ordered par

OpenStudy (tester97):

\(\sf\Huge\color{green}{Welcome~to~OpenStudy~♥}\) Now lets begin do you know how to start off solving this?

OpenStudy (anonymous):

no

OpenStudy (tester97):

Ok do you know how to change this into y intercept form?

OpenStudy (anonymous):

y=mx+b

OpenStudy (tester97):

DING DING DING :)

OpenStudy (tester97):

So we have these two equations that we need to change to y intercept form \[2x+7y=-14 \] and \[4x-3y=-19\] Do you know how to change those into y intercept form?

OpenStudy (texaschic101):

why are you changing them to y intercept form ?

OpenStudy (tester97):

Because then that way you can put it in a graphing calculator and see where the two lines intersect and get your ordered pair. Would that be right?

OpenStudy (texaschic101):

or you could just solve by elimination

OpenStudy (texaschic101):

I suppose you could do it your way...but that is not really teaching them to solve it by theirselves

OpenStudy (tester97):

True true, but everyone is different no worries though. |^_^|